Solve by the elimination method:
5r-8s=-26
8r+5s=83

Found 2 solutions by stanbon, NancyLam:
Answer by stanbon(75887)

Solve by the elimination method:
5r-8s=-26
8r+5s=83
----
Multiply thru 1st equation by 8.
Multiply thru 2nd equation by 5.
------------------------------------
40r-64s = -8*26
40r+25s = 5*83
----
Subtract the top equation from the bottome and solve for "s":
89s = 623
s = 7
---
Solve for "r":
5r-8*7 = -26
5r = -26 + 56
5r = 30
r = 6
=================
Cheers,
Stan H.

Answer by NancyLam(35)

5r-8s=-26 (mitiply 5)= 25r-40s=-130(Write equation)
8r+5s=83 (mutiply 8)= 64r+40s= 664(Write equation)

25r-40s=-130(rewrite equation)
64r+40s= 664(Write equation)
89r+0s = 534(add 2 equation)
r= 534/89 (solve for y)
r=6
subtitution r=6 for find y
5(6)-8s=-26
30 -8s =-26
s = -26-30/-8
s = 7